OpenStudy (anonymous):

Simplify the expression completely. log1/2 8 (1/2 is the base) Thanks :)

OpenStudy (anonymous):

\[\left(\frac{1}{2}\right)^x=8\] solve for \(x\)

OpenStudy (anonymous):

since \((\frac{1}{2})^3=\frac{1}{8}\) and therefore \((\frac{1}{2})^{-3}=8\) the answer is \(-3\)

OpenStudy (anonymous):

Where did you get the 3 from?

OpenStudy (anonymous):

\(2^3=8\)

OpenStudy (anonymous):

and so \((\frac{1}{2})^{-3}=8\) as well